To determine how much calcium chloride has been used, you take a sample of slush from the road to analyze. 

You collect the following data: 


    Mass of empty container: 54.54 g 

    Mass of container and sample of slush: 160.65 g 


Then, you evaporate the water from the solution and determine the mass of the residue to be 3.29 g. What is the percent by mass of CaClin the slush that was collected? 


1
Expert's answer
2021-07-21T04:40:16-0400

Given mass of CaCO3=100mg=0.1 g (Since 1g=1000mg)


molar mass of CaCO3=100 g/mol


Therefore moles of CaCO3=mass/molar mass=0.1g/100 g/mol=1x10-3 mol.


Let us take volume of solution=1L


therefore molarity of CaCO3=mol/ volume=(1x10-3 mol)/1L=1x10-3 mol/L=1x10-3 M.


The concentration or molarity of CaCO3=1x10-3 M =1 mM ( Since 1M=1000 mM)


We know that 1ppm=1mg in 1L


Here 100 mg in 1L so the concentration of CaCO3=100 ppm. (Parts-per-million).


Here I have taken volume as 1L because you didi not mension anything about volume of solution. If you provide correct value I will recalculate and give you.Thanks an di hope you like it.